# Moments of characteristic polynomials and their derivatives for $SO(2N)$ and $USp(2N)$ and their application to one-level density in families of elliptic curve $L$-functions

### v1

@@ Abstract
Using the ratios theorems, we calculate the leading order terms in $N$ for the following averages of the characteristic polynomial and its derivative: $\left|\Lambda_{A}(1)\right|^{r}\frac{\Lambda_{A}^{\prime}(\mathrm{e}^{\mathrm{i}\phi})}{\Lambda_{A}(\mathrm{e}^{\mathrm{i}\phi})}\right>_{SO(2N)}$ and $\left|\Lambda_{A}(1)\right|^{r}\frac{\Lambda_{A}^{\prime}(\mathrm{e}^{\mathrm{i}\phi})}{\Lambda_{A}(\mathrm{e}^{\mathrm{i}\phi})}\right>_{USp(2N)}$. Our expression, derived for integer $r$, permits analytic continuation in $r$ and we conjecture that this agrees with the above averages for non-integer exponents. We use this result to obtain an expression for the one level density of the ‘excised ensemble’, a subensemble of $SO(2N)$, to next-to-leading order in $N$. We then present the analogous calculation for the one level density of quadratic twists of elliptic curve $L$-functions, taking into account a number theoretical bound on the central values of the $L$-functions. The method we use to calculate the above random matrix averages uses the contour integral form of the ratios theorems, which are a key tool in the growing literature on averages of characteristic polynomials and their derivatives, and as we evaluate the next-to-leading term for large matrix size $N$, this leads to some multi-dimensional contour integrals that are slightly asymmetric in the integration variables, which might be useful in other work.
@@ 1 Introduction
[…]
Here, we calculate similar quantities over $SO(2N)$ and $USp(2N)$, with the distinction that for the purposes of our application to the one level density of zeros of $L$-functions, we do not always evaluate the characteristic polynomial at the point 1, as is the case in the aforementioned literature. To start with, in Section 2, we shall focus on the expected value
$\left|\Lambda_{A}(1)\right|^{r}\frac{\Lambda^{\prime}(\mathrm{e}^{\mathrm{i}\phi})}{\Lambda(\mathrm{e}^{\mathrm{i}\phi})}\right>_{SO(2N)},$ (2)
@@
An analogous quantity, a joint average of the Riemann zeta function and it’s logarithmic derivative, have been studied in the number theory literature for a slightly different purpose by Fazzari in [21]. Similar moments on the random matrix theory side have been considered by Simm and Wei [35].
In random matrix theory, the one level density gives us information about how the density of eigenvalues varies around the unit circle - about the likelihood of finding an eigenvalue in a particular interval on the unit circle if we draw a matrix at random from our ensemble. If one could write $\left|\Lambda_{A}(1)\right|^{r}\frac{\Lambda^{\prime}(\mathrm{e}^{\mathrm{i}\phi})}{\Lambda(\mathrm{e}^{\mathrm{i}\phi})}\right>_{SO(2N)}$ as an analytic expression in $r$, then we could use this quantity to calculate the one level density for the excised ensemble $\{A|A\in SO(2N),\Lambda_{A}(1)>\mathrm{e}^{\chi}\}$, which was introduced in [20] as a model for quadratic twist families of elliptic curve $L$-functions. In Section 5 we make the assumption that our expression for (2) holds for complex $r$ with ${\rm Re}(r)>0$.
@@ 1.2 RMT and number theory
$\int_{SO(2N)}\frac{\Lambda(1)^{K-1}\Lambda(\mathrm{e}^{-\alpha})}{\Lambda(\mathrm{e}^{-\gamma})}\mathrm{dA}$
$=$ $\frac{(-1)^{K(K-1)/2}2^{K}}{(2\pi\mathrm{i})^{K}K!}\mathrm{e}^{-N\alpha}$
[…]
@@
$\displaystyle\sum_{\begin{subarray}{c}0<d\leq X\\ \omega_{E}\chi_{d}(-M)=1\end{subarray}}\left[\frac{L_{E}^{K-1}(1/2,\chi_{d})L_{E}(1/2+\alpha,\chi_{d})}{L_{E}(1/2+\gamma,\chi_{d})}\right]$
$\qquad=\frac{(-1)^{K(K-1)/2}2^{K}}{K!(2\pi\mathrm{i})^{K})}\displaystyle\sum_{\begin{subarray}{c}0<d\leq X\\ \omega_{E}\chi_{d}(-M)=1\end{subarray}}\left(\frac{M|d|^{2}}{4\pi^{2}}\right)^{\left(-\alpha/2\right)}$
×$\times\oint\dots\oint\left(\frac{M|d|^{2}}{4\pi^{2}}\right)^{\tfrac{1}{2}\sum_{k=1}^{K}\left(w_{k}\right)}$ $\left(\frac{H_{\alpha,\gamma}(w_{1},\dots,w_{K})\Delta(w_{1}^{2},\dots,w_{K}^{2})^{2}\prod_{i=1}^{K}w_{i}\mathrm{d}w_{i}}{\displaystyle\prod_{n=1}^{K}(w_{n}-\alpha)(w_{n}+\alpha)w_{n}^{2K-2}}\right)$
@@
Here the contours enclose zero and ±$\pm\alpha$. The components such as $z(x)$ and $H_{\alpha,\gamma}$ will be discussed when we come to use these formulae in later sections (see (2) and (162) plus (6)), but the point here is to notice that if $N$ is equated to $\tfrac{1}{2}\log(M|d|^{2}/(4\pi^{2}))$, then the structure of these two formula, one for an average over matrices from $SO(2N)$ with Haar measure, and one an average over a family of elliptic curve $L$-functions, is near identical.
@@ 2 A mixed moment for $SO(2N)$
[…]
where $z(x)=(1-\mathrm{e}^{-x})^{-1}=\left(\frac{1}{x}+\frac{1}{2}+\mathcal{O}\left(x\right)\right)$, and the contours of integration enclose 0, $\alpha$ and $-\alpha$. To avoid ambiguity, we shall take the contours to be nested, i.e. the $w_{i}$ contour shall completely enclose the $w_{j}$ contour whenever $j>i$. The Vandermonde determinant is
@@
×$\times\displaystyle\prod_{n=1}^{K}\left(\frac{\exp\left(Nw_{n}\right)\left(w_{n}\right)^{3-2K}\mathrm{d}w_{n}}{z\left(w_{n}+\gamma\right)\left(w_{n}-\alpha\right)\left(w_{n}+\alpha\right)}\right).$ (15)
Note that we do not get poles from the $z(w_{i}+w_{j})$ term due to cancellation with the Vandermonde $\Delta(w_{1}^{2},\dots,w_{K}^{2})$. Each contour has poles at 00, $\alpha$ and $-\alpha$. We contract each contour onto the poles so that it consists of a small circle around each pole, with narrow necks connecting the circles. The integral will cancel on each side of the neck as there are no branch points (note we are still assuming $K$ is an integer at this point). This leaves a sum in each variable over three circular contours:
@@
×$\times\prod_{n=1}^{K}\left(\frac{\exp\left(Nw_{n}\right)\left(w_{n}\right)^{3-2K}}{z\left(w_{n}+\gamma\right)\left(w_{n}-\alpha\right)\left(w_{n}+\alpha\right)}\mathrm{d}w_{n}\right),$ (16)
@@ 2.1 Evaluating the contributions to the sum
We are aiming to evaluate $\mathcal{R}_{SO(2N)}(\alpha,\gamma,K)$ for large values of $N$. We will look at this asymptotically and also at the next-to-leading order term. To perform the sum over $\epsilon$ in (2) we need to consider separately terms where different numbers of variables are integrated round each of the three potential poles: 00, $\alpha$ and $-\alpha$. We show in the next section that in the case of a term where two or more of the $K$ variables are integrated around a pole away from 0, that term contributes zero to the sum for any value of $N$. If all variables are on contours around zero, that term will be zero at leading order (see Section 2.1.2). This leaves us with evaluating the contribution when all variables except for one are evaluated around the pole at zero, which will be considered in Section 2.1.3.
2.1.1 Two or more variables integrated around ±$\pm\alpha$
Without loss of generality, we can choose $w_{1}$ and $w_{2}$ to be integration variables which we shall be integrating around either $\alpha$ or $-\alpha$. Consider the case where $\epsilon_{1}=\epsilon_{2}=\alpha$. The only singular parts of the integrand are the factors of $(w_{1}-\alpha)$ and $(w_{2}-\alpha)$ in the denominator. Considering performing, say, the $w_{1}$ integral first, after evaluation of the residue at $w_{1}=\alpha$, this will leave us with factors of $(w_{2}-\alpha)$ from the Vandermonde in the numerator. These exactly cancel out the $(w_{2}-\alpha)$ in the denominator, so the integrand has no singularities and therefore no residues in $w_{1}$ or $w_{2}$. The same thing happens for $w_{1}=w_{2}=-\alpha$.
There are only two possible non-zero values for the $\epsilon_{i}$s, namely ±$\pm\alpha$, so if we have more than two non-zero $\epsilon_{i}$s then at least two of them must be equal and the integral will vanish.
Let us next consider the case where two of the $\epsilon_{i}$s are non-zero and non-equal. For the purposes of doing the integrals in $w_{1}$ and $w_{2}$, the other $w_{i}$s may be treated as constants. Let us consider only the parts of the integrand that either diverge or tend to zero as $w_{1}$ and $w_{2}$ approach ±$\pm\alpha$. We shall show that this vanishes when we do the contour integrals in $w_{1}$ and $w_{2}$. This may be achieved by showing that
@@ 2.1.1 Two or more variables integrated around ±$\pm\alpha$
This has no poles in $w_{2}$, since the $(\alpha+w_{2})^{-1}$ term from the series expansion of $z(\alpha+w_{2})$ is cancelled out. This completes our demonstration that terms in (2) with two or more residues at evaluated at ±$\pm\alpha$ will have zero contribution to the total sum.
@@ 2.1.2 All the poles at zero
We have established that terms with fewer than $K-1$ poles at zero do not contribute at all to the sum for any value of $N$. Let us now evaluate the contribution from the terms where all of the poles are integrated around zero, but this time we will scale the variables of integration by $1/N$ and look at the leading order term. We will find that even the leading order term of this calculation is negligible compared to the size of other terms in the sum (2) (see Section 2.1.3), but we can show that the coefficient of the leading order term here is zero, and as the method is useful later, we will demonstrate it in full detail. We start with the terms in (2) where all poles are evaluated at zero:
@@
×$\times\prod_{n=1}^{K}\left(\frac{\left(\frac{w_{n}}{N}\right)^{3-2K}\mathrm{d}\left(\frac{w_{n}}{N}\right)}{z\left(\frac{w_{n}}{N}+\gamma\right)\left(\frac{w_{n}}{N}-\alpha\right)\left(\frac{w_{n}}{N}+\alpha\right)}\right).$ (19)
@@
×$\times\prod_{n=1}^{K}\left(\frac{w_{n}}{N}\right)^{3-2K}\mathrm{d}\left(\frac{w_{n}}{N}\right)\;\left(1+O(1/N)\right),$ (20)
@@
×$\qquad\times\oint_{0}\dots\oint_{0}\displaystyle\prod_{1\leq j<k\leq K}\left(w_{j}-w_{k}\right)\left(w_{j}^{2}-w_{k}^{2}\right)\exp\left(\displaystyle\sum_{l=1}^{K}w_{l}\right)\prod_{n=1}^{K}w_{n}^{3-2K}\mathrm{d}w_{n}\;\left(1+O(1/N)\right).$
@@
where $S_{n}$ is the permutation group of size $n$, where for convenience we have permutations of the numbers $10,1,2,\;\ldots,n-$ 1. Using (22), we can rewrite the Vandermonde in terms of sums over the permutation group of size $K$:
$\frac{C^{SO(2N)}_{\alpha,\gamma,K}N^{\tfrac{1}{2}K^{2}-\tfrac{5}{2}K}}{z(\gamma)^{K}(-\alpha^{2})^{K}}\oint_{0}\dots\oint_{0}\exp\left(\displaystyle\sum_{l=1}^{K-1}w_{l}\right)\left(\displaystyle\sum_{\sigma\in S_{K}}\mathrm{sgn}(\sigma)w_{1}^{2\sigma_{0}}w_{2}^{2\sigma_{1}}\dots w_{K}^{2\sigma_{K-1}}\right)$
×$\qquad\times\left(\displaystyle\sum_{\rho\in S_{K}}\mathrm{sgn}(\rho)w_{1}^{\rho_{0}}w_{2}^{\rho_{1}}\dots w_{K}^{\rho_{K-1}}\right)\prod_{n=1}^{K}w_{n}^{3-2K}\mathrm{d}w_{n}\;\left(1+O(1/N)\right).$
In each term of the sum over $\rho\in S_{K}$ in (2.1.2), let us relabel the variables of integration so that $w_{1}$ appears with exponent 0, $w_{2}$ with exponent 1 etc. This results in $K!$ identical terms, and the sign change from the relabelling cancels with the sign in the sum over $\sigma$. So, at leading order we have
@@
×$\times\left(\displaystyle\sum_{\sigma\in S_{K}}\mathrm{sgn}(\sigma)w_{1}^{2\sigma_{0}}w_{2}^{2\sigma_{1}}\dots w_{K}^{2\sigma_{K-1}}\right)w_{1}^{3-2K}w_{2}^{4-2K}\dots w_{K}^{2-K}\mathrm{d}w_{1}\dots\mathrm{d}w_{K}$
@@
×$\times\left|\begin{array}[]{ccccc}1&w_{1}^{2}&w_{1}^{4}&\dots&w_{1}^{2K-2}\\ 1&w_{2}^{2}&w_{2}^{4}&\dots&w_{2}^{2K-2}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1&w_{K}^{2}&w_{K}^{4}&\dots&w_{K}^{2K-2}\end{array}\right|w_{1}^{3-2K}w_{2}^{4-2K}\dots w_{K}^{2-K}\mathrm{d}w_{1}\dots\mathrm{d}w_{K}$ (28)
@@
×$\times\left|\begin{array}[]{ccccc}\oint\mathrm{e}^{w_{1}}w_{1}^{3-2K}\mathrm{d}w_{1}&\oint\mathrm{e}^{w_{1}}w_{1}^{5-2K}\mathrm{d}w_{1}&\dots&\oint\mathrm{e}^{w_{1}}w_{1}^{-1}\mathrm{d}w_{1}&\oint\mathrm{e}^{w_{1}}w_{1}^{1}\mathrm{d}w_{1}\\ \oint\mathrm{e}^{w_{2}}w_{2}^{4-2K}\mathrm{d}w_{2}&\oint\mathrm{e}^{w_{2}}w_{2}^{6-2K}\mathrm{d}w_{2}&\dots&\oint\mathrm{e}^{w_{2}}w_{2}^{0}\mathrm{d}w_{2}&\oint\mathrm{e}^{w_{1}}w_{1}^{2}\mathrm{d}w_{1}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \oint\mathrm{e}^{w_{K}}w_{K}^{2-K}\mathrm{d}w_{K}&\oint\mathrm{e}^{w_{K}}w_{K}^{4-K}\mathrm{d}w_{K}&\dots&\oint\mathrm{e}^{w_{K}}w_{K}^{K-2}\mathrm{d}w_{K}&\oint\mathrm{e}^{w_{K}}w_{K}^{K}\mathrm{d}w_{K}\end{array}\right|.$ (33)
@@
where the contour of integration comes in from +$\infty$+ ∞, encircles the origin in the anticlockwise direction and then returns to +$\infty$+ ∞. The exponent on $t$ will always be an integer in our case, so the integral on the two strands to +$\infty$+ ∞ will cancel, leaving just an anticlockwise circle around the origin. So, (2.1.2) becomes
@@
and this is the leading order contribution from the term in (2) where all residues are evaluated at zero. Since $01/\Gamma(n)=$ 0 when $n$ is a negative integer, the final column consists entirely of zeros, which means the determinant is zero.

### v2

@@ Abstract
Using the ratios theorems, we calculate the leading order terms in $N$ for the following averages of the characteristic polynomial and its derivative: $\left<\left|\Lambda_{A}(1)\right|^{r}\frac{\Lambda_{A}^{\prime}(\mathrm{e}^{\mathrm{i}\phi})}{\Lambda_{A}(\mathrm{e}^{\mathrm{i}\phi})}\right>_{SO(2N)}$ and $\left<\left|\Lambda_{A}(1)\right|^{r}\frac{\Lambda_{A}^{\prime}(\mathrm{e}^{\mathrm{i}\phi})}{\Lambda_{A}(\mathrm{e}^{\mathrm{i}\phi})}\right>_{USp(2N)}$. Our expression, derived for integer $r$, permits analytic continuation in $r$ and we conjecture that this agrees with the above averages for non-integer exponents. We use this result to obtain an expression for the one level density of the ‘excised ensemble’, a subensemble of $SO(2N)$, to next-to-leading order in $N$. We then present the analogous calculation for the one level density of quadratic twists of elliptic curve $L$-functions, taking into account a number theoretical bound on the central values of the $L$-functions. The method we use to calculate the above random matrix averages uses the contour integral form of the ratios theorems, which are a key tool in the growing literature on averages of characteristic polynomials and their derivatives, and as we evaluate the next-to-leading term for large matrix size $N$, this leads to some interesting multi-dimensional contour integrals, defined at the end of Section 2.1.3, which might be useful in other work.
@@ 1 Introduction
Mixed moments for $U(N)$, in the spirit of $\left<\left|\Lambda_{A}(1)\right|^{k}\left|\Lambda_{A}^{\prime}(1)\right|^{r}\right>_{A\in U(N)}$ (actually in some cases the results are for a variation of the characteristic polynomial that is multiplied by a factor so as to be real on the unit circle) have been studied extensively, starting with Hughes in 2001 [27], by Dehaye [21, 22], by Winn [40] and Conrey, Rubinstein and Snaith in [16], and more recently in two large collaborations [8] and [11]. These results have led to an understanding of the rate of growth of the leading order term when the matrix size $N$ is large, and several varied expressions for the leading order coefficient in terms of determinants, combinatorial sums, integrals or relating it to solutions of Painlevé equations. This has been extended to circular beta ensembles by Forrester [25] and to higher derivatives in $U(N)$ by Keating and Wei [32, 33], Barhoumi-Andréani [9] and Assiotis, Gunes, Keating and Wei [5] and generalised to $USp(2N)$, $SO(2N)$ and $O^{-}(2N)$ by Altuǧ et al [1] and Andrade and Best [3]. There is an ongoing effort to give explicit expressions for the leading order coefficient also when the exponents $k$ and $r$ are non-integer. Early results in this direction generalised to $k\in\mathbb{Z},\ 2r\in\mathbb{Z}$ [40] or $k\in\mathbb{R}$, $r\in\mathbb{Z}$ [22], and in the past few years Assiotis, Keating and Warren [7] have extended both the exponent on the characteristic polynomial and its derivative to real numbers and expressed the leading order coefficient as the expectation value of a certain random variable, with an extremely recent extension of this to averages of products of many derivatives of any order raised to any positive real powers in [5]. Similar work has been carried out for other ensembles by Assiotis, Gunes and Soor in [6].
Here, we calculate similar quantities over $SO(2N)$ and $USp(2N)$, with the distinction that for the purposes of our application to the one level density of zeros of $L$-functions, we do not always evaluate the characteristic polynomial at the point 1, as is the case in the aforementioned literature. In this way it is more akin to the $n$-correlation and $n$-level density calculations in [17, 19, 18, 35]. To start with, in Section 2, we shall focus on the expected value
$\left<\left|\Lambda_{A}(1)\right|^{r}\frac{\Lambda^{\prime}(\mathrm{e}^{\mathrm{i}\phi})}{\Lambda(\mathrm{e}^{\mathrm{i}\phi})}\right>_{SO(2N)},$ (2)
@@
An analogous quantity, a joint average of the Riemann zeta function and its logarithmic derivative, have been studied in the number theory literature for a slightly different purpose by Fazzari in [24]. Similar moments on the random matrix theory side have been considered by Simm and Wei [38].
In random matrix theory, the one level density gives us information about how the density of eigenvalues varies around the unit circle - about the likelihood of finding an eigenvalue in a particular interval on the unit circle if we draw a matrix at random from our ensemble. If one could write $\left<\left|\Lambda_{A}(1)\right|^{r}\frac{\Lambda^{\prime}(\mathrm{e}^{\mathrm{i}\phi})}{\Lambda(\mathrm{e}^{\mathrm{i}\phi})}\right>_{SO(2N)}$ as an analytic expression in $r$, then we could use this quantity to calculate the one level density for the excised ensemble $\{A|A\in SO(2N),\Lambda_{A}(1)>\mathrm{e}^{\chi}\}$, which was introduced in [23] as a model for quadratic twist families of elliptic curve $L$-functions. In Section 5 we make the assumption that our expression for (2) holds for complex $r$ with ${\rm Re}(r)>0$.
@@ 1.2 RMT and number theory
$\int_{SO(2N)}\frac{\Lambda(1)^{K-1}\Lambda(\mathrm{e}^{-\alpha})}{\Lambda(\mathrm{e}^{-\gamma})}\mathrm{dA}=\frac{(-1)^{K(K-1)/2}2^{K}}{(2\pi\mathrm{i})^{K}K!}\mathrm{e}^{-N\alpha}$
$\qquad\times\oint\dots\oint\exp\left(N\displaystyle\sum_{l=1}^{K}w_{l}\right)\frac{H^{SO(2N)}_{\alpha,\gamma}(w_{1},\dots,w_{K})\Delta(w_{1}^{2},\dots,w_{K}^{2})^{2}\prod_{i=1}^{K}w_{i}\mathrm{d}w_{i}}{\displaystyle\prod_{n=1}^{K}(w_{n}-\alpha)(w_{n}+\alpha)w_{n}^{2K-2}},$ (8)
@@
$\displaystyle\sum_{\begin{subarray}{c}0<d\leq X\\ \omega_{E}\chi_{d}(-M)=1\end{subarray}}\left[\frac{L_{E}^{K-1}(1/2,\chi_{d})L_{E}(1/2+\alpha,\chi_{d})}{L_{E}(1/2+\gamma,\chi_{d})}\right]=\frac{(-1)^{K(K-1)/2}2^{K}}{K!(2\pi\mathrm{i})^{K})}\displaystyle\sum_{\begin{subarray}{c}0<d\leq X\\ \omega_{E}\chi_{d}(-M)=1\end{subarray}}\left(\frac{M|d|^{2}}{4\pi^{2}}\right)^{\left(-\alpha/2\right)}$
$\times\oint\dots\oint\left(\frac{M|d|^{2}}{4\pi^{2}}\right)^{\tfrac{1}{2}\sum_{k=1}^{K}\left(w_{k}\right)}\left(\frac{H_{\alpha,\gamma}(w_{1},\dots,w_{K})\Delta(w_{1}^{2},\dots,w_{K}^{2})^{2}\prod_{i=1}^{K}w_{i}\mathrm{d}w_{i}}{\displaystyle\prod_{n=1}^{K}(w_{n}-\alpha)(w_{n}+\alpha)w_{n}^{2K-2}}\right)$
@@
Here the contours enclose zero and $\pm\alpha$. The components such as $z(x)$ and $H_{\alpha,\gamma}$ will be discussed when we come to use these formulae in later sections (see (2) and (162) plus (6)), but the point here is to notice that if $N$ is equated to $\tfrac{1}{2}\log(M|d|^{2}/(4\pi^{2}))$, then the structure of these two formula, one for an average over matrices from $SO(2N)$ with Haar measure, and one an average over a family of elliptic curve $L$-functions, is near identical.
@@ 2 A mixed moment for $SO(2N)$
$\times\oint\dots\oint$ $\frac{\displaystyle\prod_{1\leq j<k\leq K}z(w_{j}+w_{k})z(2\gamma)\exp\left(N\displaystyle\sum_{l=1}^{K}w_{l}\right)\Delta(w_{1}^{2},\dots,w_{K}^{2})^{2}\prod_{i=1}^{K}w_{i}\mathrm{d}w_{i}}{\displaystyle\prod_{n=1}^{K}z(w_{n}+\gamma)(w_{n}-\alpha)(w_{n}+\alpha)w_{n}^{2K-2}},$ (12)
where $z(x)=(1-\mathrm{e}^{-x})^{-1}=\left(\frac{1}{x}+\frac{1}{2}+\mathcal{O}\left(x\right)\right)$, ${\rm Re}(\gamma)>0$, and the contours of integration enclose 0, $\alpha$ and $\alpha$. To avoid ambiguity, we shall take the contours to be nested, i.e. the $w_{i}$ contour shall completely enclose the $w_{j}$ contour whenever $j>i$. The Vandermonde determinant is
@@
$\times\displaystyle\prod_{n=1}^{K}\left(\frac{\exp\left(Nw_{n}\right)\left(w_{n}\right)^{3-2K}\mathrm{d}w_{n}}{z\left(w_{n}+\gamma\right)\left(w_{n}-\alpha\right)\left(w_{n}+\alpha\right)}\right).$ (15)
Note that we do not get poles from the $z(w_{i}+w_{j})$ term due to cancellation with the Vandermonde $\Delta(w_{1}^{2},\dots,w_{K}^{2})$. Each contour has poles at 0, $\alpha$ and $\alpha$. We contract each contour onto the poles so that it consists of a small circle around each pole, with narrow necks connecting the circles. The integral will cancel on each side of the neck as there are no branch points (note we are still assuming $K$ is an integer at this point). This leaves a sum in each variable over three circular contours:
@@
$\times\prod_{n=1}^{K}\left(\frac{\exp\left(Nw_{n}\right)\left(w_{n}\right)^{3-2K}}{z\left(w_{n}+\gamma\right)\left(w_{n}-\alpha\right)\left(w_{n}+\alpha\right)}\mathrm{d}w_{n}\right),$ (16)
@@ 2.1 Evaluating the contributions to the sum
We are aiming to evaluate $\mathcal{R}_{SO(2N)}(\alpha,\gamma,K)$ for large values of $N$. We will look at this asymptotically and also at the next-to-leading order term. To perform the sum over $\epsilon$ in (2) we need to consider separately terms where different numbers of variables are integrated round each of the three potential poles: 0, $\alpha$ and $\alpha$. We show in the next section that in the case of a term where two or more of the $K$ variables are integrated around a pole away from 0, that term contributes zero to the sum for any value of $N$. If all variables are on contours around zero, that term will be zero at leading order (see Section 2.1.2). This leaves us with evaluating the contribution when all variables except for one are evaluated around the pole at zero, which will be considered in Section 2.1.3.
2.1.1 Two or more variables integrated around $\pm\alpha$
Without loss of generality, we can choose $w_{1}$ and $w_{2}$ to be integration variables which we shall be integrating around either $\alpha$ or $\alpha$. Consider the case where $\epsilon_{1}=\epsilon_{2}=\alpha$. The only singular parts of the integrand are the factors of $(w_{1}-\alpha)$ and $(w_{2}-\alpha)$ in the denominator. Considering performing, say, the $w_{1}$ integral first, after evaluation of the residue at $w_{1}=\alpha$, this will leave us with factors of $(w_{2}-\alpha)$ from the Vandermonde in the numerator. These exactly cancel out the $(w_{2}-\alpha)$ in the denominator, so the integrand has no singularities and therefore no residues in $w_{1}$ or $w_{2}$. The same thing happens for $w_{1}=w_{2}=-\alpha$.
There are only two possible non-zero values for the $\epsilon_{i}$s, namely $\pm\alpha$, so if we have more than two non-zero $\epsilon_{i}$s then at least two of them must be equal and the integral will vanish.
Let us next consider the case where two of the $\epsilon_{i}$s are non-zero and non-equal. For the purposes of doing the integrals in $w_{1}$ and $w_{2}$, the other $w_{i}$s may be treated as constants. Let us consider only the parts of the integrand that either diverge or tend to zero as $w_{1}$ and $w_{2}$ approach $\pm\alpha$. We shall show that this vanishes when we do the contour integrals in $w_{1}$ and $w_{2}$. This may be achieved by showing that
@@ 2.1.1 Two or more variables integrated around $\pm\alpha$
This has no poles in $w_{2}$, since the $(\alpha+w_{2})^{-1}$ term from the series expansion of $z(\alpha+w_{2})$ is cancelled out. This completes our demonstration that terms in (2) with two or more residues at evaluated at $\pm\alpha$ will have zero contribution to the total sum.
@@ 2.1.2 All the poles at zero
We have established that terms with fewer than K1K-1 poles at zero do not contribute at all to the sum for any value of $N$. Let us now evaluate the contribution from the terms where all of the poles are integrated around zero, but this time we will scale the variables of integration by $1/N$ and look at the leading order term. We will find that even the leading order term of this calculation is negligible compared to the size of other terms in the sum (2) (see Section 2.1.3), but we can show that the coefficient of the leading order term here is zero, and as the method is useful later, we will demonstrate it in full detail. We start with the terms in (2) where all poles are evaluated at zero:
@@
$\times\prod_{n=1}^{K}\left(\frac{\left(\frac{w_{n}}{N}\right)^{3-2K}\mathrm{d}\left(\frac{w_{n}}{N}\right)}{z\left(\frac{w_{n}}{N}+\gamma\right)\left(\frac{w_{n}}{N}-\alpha\right)\left(\frac{w_{n}}{N}+\alpha\right)}\right).$ (19)
@@
$\times\prod_{n=1}^{K}\left(\frac{w_{n}}{N}\right)^{3-2K}\mathrm{d}\left(\frac{w_{n}}{N}\right)\;\left(1+O(1/N)\right),$ (20)
@@
$\qquad\times\oint_{0}\dots\oint_{0}\displaystyle\prod_{1\leq j<k\leq K}\left(w_{j}-w_{k}\right)\left(w_{j}^{2}-w_{k}^{2}\right)\exp\left(\displaystyle\sum_{l=1}^{K}w_{l}\right)\prod_{n=1}^{K}w_{n}^{3-2K}\mathrm{d}w_{n}\;\left(1+O(1/N)\right).$
@@
where $S_{n}$ is the permutation group of size $n$, where for convenience we have permutations of the numbers $0,1,2,\;\ldots,n-1$. Using (22), we can rewrite the Vandermonde in terms of sums over the permutation group of size $K$:
$\frac{C^{SO(2N)}_{\alpha,\gamma,K}N^{\tfrac{1}{2}K^{2}-\tfrac{5}{2}K}}{z(\gamma)^{K}(-\alpha^{2})^{K}}\oint_{0}\dots\oint_{0}\exp\left(\displaystyle\sum_{l=1}^{K-1}w_{l}\right)\left(\displaystyle\sum_{\sigma\in S_{K}}\mathrm{sgn}(\sigma)w_{1}^{2\sigma_{0}}w_{2}^{2\sigma_{1}}\dots w_{K}^{2\sigma_{K-1}}\right)$ (23)
$\qquad\times\left(\displaystyle\sum_{\rho\in S_{K}}\mathrm{sgn}(\rho)w_{1}^{\rho_{0}}w_{2}^{\rho_{1}}\dots w_{K}^{\rho_{K-1}}\right)\prod_{n=1}^{K}w_{n}^{3-2K}\mathrm{d}w_{n}\;\left(1+O(1/N)\right).$
In each term of the sum over $\rho\in S_{K}$ in (23), let us relabel the variables of integration so that $w_{1}$ appears with exponent 0, $w_{2}$ with exponent 1 etc. This results in $K!$ identical terms, and the sign change from the relabelling cancels with the sign in the sum over $\sigma$. So, at leading order we have
@@
$\times\left(\displaystyle\sum_{\sigma\in S_{K}}\mathrm{sgn}(\sigma)w_{1}^{2\sigma_{0}}w_{2}^{2\sigma_{1}}\dots w_{K}^{2\sigma_{K-1}}\right)w_{1}^{3-2K}w_{2}^{4-2K}\dots w_{K}^{2-K}\mathrm{d}w_{1}\dots\mathrm{d}w_{K}$
@@
$\times\left|\begin{array}[]{ccccc}1&w_{1}^{2}&w_{1}^{4}&\dots&w_{1}^{2K-2}\\ 1&w_{2}^{2}&w_{2}^{4}&\dots&w_{2}^{2K-2}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1&w_{K}^{2}&w_{K}^{4}&\dots&w_{K}^{2K-2}\end{array}\right|w_{1}^{3-2K}w_{2}^{4-2K}\dots w_{K}^{2-K}\mathrm{d}w_{1}\dots\mathrm{d}w_{K}$ (28)
@@
$\times\left|\begin{array}[]{ccccc}\oint\mathrm{e}^{w_{1}}w_{1}^{3-2K}\mathrm{d}w_{1}&\oint\mathrm{e}^{w_{1}}w_{1}^{5-2K}\mathrm{d}w_{1}&\dots&\oint\mathrm{e}^{w_{1}}w_{1}^{-1}\mathrm{d}w_{1}&\oint\mathrm{e}^{w_{1}}w_{1}^{1}\mathrm{d}w_{1}\\ \oint\mathrm{e}^{w_{2}}w_{2}^{4-2K}\mathrm{d}w_{2}&\oint\mathrm{e}^{w_{2}}w_{2}^{6-2K}\mathrm{d}w_{2}&\dots&\oint\mathrm{e}^{w_{2}}w_{2}^{0}\mathrm{d}w_{2}&\oint\mathrm{e}^{w_{1}}w_{1}^{2}\mathrm{d}w_{1}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \oint\mathrm{e}^{w_{K}}w_{K}^{2-K}\mathrm{d}w_{K}&\oint\mathrm{e}^{w_{K}}w_{K}^{4-K}\mathrm{d}w_{K}&\dots&\oint\mathrm{e}^{w_{K}}w_{K}^{K-2}\mathrm{d}w_{K}&\oint\mathrm{e}^{w_{K}}w_{K}^{K}\mathrm{d}w_{K}\end{array}\right|.$ (33)
@@
where the contour of integration comes in from $\infty$, encircles the origin in the anticlockwise direction and then returns to $\infty$. The exponent on $t$ will always be an integer in our case, so the integral on the two strands to $\infty$ will cancel, leaving just an anticlockwise circle around the origin. So, (2.1.2) becomes
@@
and this is the leading order contribution from the term in (2) where all residues are evaluated at zero. Since $1/\Gamma(n)=0$ when $n$ is a negative integer, the final column consists entirely of zeros, which means the determinant is zero.